In the latest Indian Jones film, Indy is supposed to throw a grenade from his car, which is going 72.0 km/h , to his enemy's car, which is going 128 km/h . The enemy's car is 14.3 m in front of the Indy's when he lets go of the grenade.

Part A

If Indy throws the grenade so its initial velocity relative to him is at an angle of 45 ∘ above the horizontal, what should the magnitude of the initial velocity be? The cars are both traveling in the same direction on a level road. You can ignore air resistance.

Part B

Find the magnitude of the velocity relative to the earth.

Hint: Use Galileo's equation where Frame A is ground, Frame B is Indy's car and the object of interest P is the grenade. Use your answer in part A for speed and given direction to calculate the x and y components of the velocity with respect to Indy's car. The velocity of Indy's car with respect to the ground is straightforward to write in x and y components.
Answer #1

Part A:

The range of projectile at an angle of 45 above the horizontal with an initial speed u will be

R-u-sin 2 x 45°

If you consider that Indi´s car is at rest and then enemy's car will go at a speed of (128 - 72) or 56 Km/h i.e. 15.56 m/s. The vertical component of speed of the grenade is u sin 45o or u/\sqrt{2} . Therefore, the time of flight will be

2u sin 45° 2u T =

Therefore, the distance traveled by the enemy´s car will be 15.56 T which must equal to the range of the projectile.

2u x 15.56 = 9

1.4 = 15.56V2 = 22 m/s = 79.2 Km/h

The magnitude of the initial velocity will be 79.2 Km/h.

Part B:

The velocity of the grenade along the horizontal and vertical direction (as the angle is 45o) are 15.56 m/s or 56 Km/h. The velocity of Indy´s car in horizontal direction is 72 Km/h. Therefore, the grenade velocity with respect to earth along horizontal direction will be (72 Km/h + 56 Km/h) or 128 Km/h. The vertical component of grenade velocity with respect to Earth will be 56 Km/h. Therefore, the resultant velocity with respect to earth will be

V = V1282 +562 = 139.7 Km/h = 38.8 m/s

the magnitude of the velocity relative to the earth will be 139.7 Km/h